# SimpleDeepSearcher: Deep Information Seeking via Web-Powered Reasoning Trajectory Synthesis

### v2

@@ Abstract
Retrieval-augmented generation (RAG) systems have advanced large language models (LLMs) in complex deep search scenarios requiring multi-step reasoning and iterative information retrieval. However, existing approaches face critical limitations that lack high-quality training trajectories or suffer from the distributional mismatches in simulated environments and prohibitive computational costs for real-world deployment. This paper introduces SimpleDeepSearcher, a lightweight yet effective framework that bridges this gap through strategic data engineering rather than complex training paradigms. Our approach synthesizes high-quality training data by simulating realistic user interactions in live web search environments, coupled with a multi-criteria curation strategy that optimizes the diversity and quality of input and output side. Experiments on five benchmarks across diverse domains demonstrate that SFT on only 871 curated samples yields significant improvements over RL-based baselines. Our work establishes SFT as a viable pathway by systematically addressing the data-scarce bottleneck, offering practical insights for efficient deep search systems. Our code is available at https://github.com/RUCAIBox/SimpleDeepSearcher.
@@
Shuang Sun1thanks:    Equal contributions., Huatong Song [^1] footnotemark: 1, Yuhao Wang1, Ruiyang Ren1, Jinhao Jiang1, Junjie Zhang1, Fei Bai1, Jia Deng1, Wayne Xin Zhao1thanks:    Corresponding author., Zheng Liu2, Lei Fang3, Zhongyuan Wang2, Ji-Rong Wen1 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence 3DataCanvas Alaya NeW {sunshuanguns, batmanfly}@gmail.com
@@ 1 Introduction
Figure 1: Overall framework of our proposed SimpleDeepSearcher approach. $p$ denotes the prompt, $r$ denotes the reasoning content, $q$ represents the search query, and $d$ refers to the retrieved document after summarization. $t_{s}$ and $t_{e}$ are special tokens indicating the beginning and end of the search query, and $a$ denotes the final answer.
In recent years, retrieval-augmented generation (RAG) methods have significantly enhanced LLMs by incorporating external knowledge retrieval mechanisms (Zhao et al., 2024; Gao et al., 2024). Recent advancements have extended these capabilities to complex deep search scenarios that demand multi-step reasoning with iterative information retrieval and synthesis(Alzubi et al., 2025).
To address the complex reasoning demands in deep search scenarios, early research explored prompt-based strategies that guide models to decompose questions, generate queries, and retrieve information iteratively (Jiang et al., 2024; Teng et al., 2025; Li et al., 2025a). Other studies have attempted to improve model performance through supervised fine-tuning (SFT) (Wang et al., 2025), but due to the lack of high-quality interaction trajectories for reasoning and retrieval, SFT methods often fail to fully activate the model’s autonomous search abilities (Schick et al., 2023). To further enhance the model’s autonomous search capabilities, reinforcement learning (RL) (Sutton et al., 1999) is considered as a promising solution to train models through real-time interaction with the environment (Song et al., 2025; Jin et al., 2025; Zheng et al., 2025). However, most RL-based approaches operate within artificial environments using static document corpora, creating a distributional mismatch with real-world web dynamics. Moreover, the inherent computational intensity of RL training escalates exponentially when interfacing with live search APIs (Sun et al., 2025).
Given the overhead and complexity of RL-based training, we posits that SFT remains a viable pathway for building efficient deep search systems. While SFT offers a streamlined training process, it faces the critical challenge of lacking high-quality training data in deep search scenarios. On the input side, existing QA datasets often lack the diversity and complexity of questions and search-oriented purposes on the Web, which are essential for deep search training. On the output side, traditional answer annotations omit the critical reasoning traces (search operations, evidence synthesis, and efficient decision paths) required for teaching search-integrated reasoning strategies.
@@ 2.1 Overview
First, we replace static document retrieval with real-time network interactions, simulating human search behavior through an iterative cycle of “reasoning-searching-summarizing-generating.” By directly processing raw HTML content via commercial search APIs, we capture diverse web information features—ranging from structured data snippets to unstructured narrative discourse. Based on this, we first filter input queries using domain heterogeneity, keyword diversity, and knowledge unit complexity to construct a maximally informative training foundation while ensuring selected queries align with real-world web search scenarios. Additionally, we apply a filtering mechanism to LLM-synthesized responses, implementing a four-dimensional quality filter that simultaneously optimizes format standardization, reasoning path control, question difficulty, and search effectiveness to guarantee response quality.
The framework’s modular design offers three distinctive advantages over conventional approaches: First, it exposes the model to authentic search artifacts and noise patterns through real web interactions. Second, our multidimensional filtering strategy enables state-of-the-art performance with remarkably small SFT datasets, eliminating dependency on resource-heavy RL training. Third, the decoupled architecture between data synthesis and model constraints provides exceptional flexibility that our curated datasets can enhance any LLMs while maintaining compatibility with emerging reasoning architectures and alternative training paradigms including RL. Since the searched content is not generated by the LLM, we mask out these tokens during the SFT process.
@@ 2.2 Data Synthesis in Real Web Environment
Typically, traditional retrieval-augmented generation (RAG) systems rely on closed and static knowledge corpora. Such knowledge corpora exhibit two primary limitations: firstly, the content they contain often consists of refined and condensed segments (Barnett et al., 2024); secondly, the information within these knowledge corpora lacks timeliness (Ouyang et al., 2025). Consequently, RAG systems are limited in their ability to simulate authentic user search behaviors, as users typically search within open, dynamic, and complex web environments where the information is not only diverse in format and varied in quality but is also frequently accompanied by redundancy and noise. In light of this, our data synthesis approach does not rely on curated document collections but is instead grounded in the real, open web environment. This authentic web environment also places greater demands on the model’s capabilities for information extraction, synthesis, and reasoning.
@@
Our data synthesis strategy is firmly rooted in real web scenarios, which substantially enriches the diversity and authenticity of training samples. Our strategy generates more practical and representative supervisory signals for SFT, thereby addressing critical limitations in conventional SFT paradigms.
@@ 2.3 Diversity-aware Query Sampling
1:Annotated dataset $D$ with domains $d_{1},d_{2},\dots,d_{m}$, target number of queries $N$
2:$N_{d}\leftarrow N/m$
3:$S\leftarrow\emptyset$triangleright Initialize the target set
@@
9:         $K\leftarrow\emptyset$triangleright Initialize the keyword set
@@
To engineer a deep search architecture with advanced query comprehension and reasoning capabilities, we implement a strategic repurposing of open-domain question answering (QA) resources. These curated datasets offer natural language questions that inherently require multi-hop information retrieval operations, thereby exhibiting strong task alignment with the cognitive demands of deep search systems. Our selection protocol combines single-hop and multi-hop QA benchmarks through principled composition, ensuring coverage of both atomic and composite reasoning paradigms.
However, empirical evidence suggests that naive dataset scaling yields diminishing returns in SFT. The efficacy of such approaches fundamentally depends on the intrinsic diversity and informational entropy of training instances. While existing open-domain QA corpora provide substantial volume, systematic analysis reveals three critical limitations: (1) domain-specific overrepresentation creating skewed knowledge distributions, (2) repetitive syntactic patterns reducing linguistic variability, and (3) semantic simplicity thresholds below real-world query complexity. These factors collectively induce model brittleness and constrain cross-domain generalization potential. To address these critical limitations, we introduce a diversity-aware query sampling strategy to implement systematic data filtering through tripartite orthogonal criteria:
Keyword Diversity focuses on the distributional diversity of core semantic constituents. we ensure non-redundant exposure to low-frequency conceptual entities, multi-order relational dependencies, and contextually ambiguous referential expressions. Such systematic variation compels the model to transcend superficial lexical pattern matching, instead developing reasoning architectures essential for interpreting complex entity interactions.
@@ 2.4 Multi-Dimention Response Curation
The presence of low-quality reasoning outputs in language models not only compromises performance and transparency but also introduces noise into training signals, leading to inefficient computational resource utilization. To address these challenges, we developed a systematic filtering protocol that selects optimal solutions through rigorous and comprehensive evaluation of multiple responses per query.
@@ 3.1 Experimental Setup
We sample training data from single-hop and multi-hop knowledge-intensive QA datasets to cover a wide range of domains and question difficulty. For single-hop questions, we use Natural Questions (Kwiatkowski et al., 2019) and SimpleQA (Wei et al., 2024). For multi-hop questions, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), MuSiQue (Tang and Yang, 2024), and MultiHopRAG (Tang and Yang, 2024). To test the model’s performance on out-of-domain data, we select Bamboogle (Press et al., 2022), FRAMES (Krishna et al., 2024), and GAIA (Mialon et al., 2023) as extra test sets. These datasets are not used during training and help evaluate how well the model works on new domains. We evaluate our approach on 500 randomly sampled instances from the validation sets of HotpotQA, 2WikiMultiHopQA, and MuSiQue. For Bamboogle and FRAMES, we use their full test sets. For GAIA, we use 103 examples from the text-only validation subset (Li et al., 2025b).
@@
We consider following type of baselines: Naive Generation: Direct generation of answers without retrieval. Standard RAG (Zhao et al., 2024): Directly retrieve relevant documents by querying the original question. Search-o1 (Li et al., 2025a): Encourages the model to perform self-initiated retrieval using prompts. RAG-RL: R1-Searcher (Song et al., 2025) and DeepResearcher (Zheng et al., 2025), the open-source 7B model trained with reinforcement learning to enable self-initiated retrieval. We conduct experiments using the following model backbones with an online search engine, including Qwen-2.5-7B-Instruct, Qwen-2.5-32B-Instruct, Deepseek-Distilled-Qwen-2.5-32B, and QwQ-32B.
@@
Our experimental setup consists of four main components: query sampling, data synthesis, generation, and SFT. During query sampling, we used QwQ-32B to annotate each query with its corresponding domain and keywords. For data synthesis, we employed QwQ-32B as the reasoning model and Google Search API as the search engine, with a maximum of 10 search calls and 15 reasoning turns per query. For each query, we sampled 10 candidate responses. For generation, all models are configured with a maximum sequence length of 20,480 tokens, temperature of 0.6, top-p of 0.95, and top-k of 40. In the SFT phase, we use a total batch size of 64 and train for 6 epochs with a learning rate of 1e-5, warmup ratio of 0.03, and a sequence length of 30,000 tokens. During fine-tuning, external retrieval documents are masked to avoid learning from noisy or spurious information. All prompts used in the experiments are provided in Appendix E.
3.2 Main Results
@@ 4.1 Ablation Study
To validate the effectiveness of the proposed SimpleDeepSearcher, we conduct a comprehensive ablation analysis using Qwen2.5-7B-Instruct on the Bamboogle and GAIA datasets. We conduct detailed ablation studies on three main aspects: (1) Query Sampling: w/o DH removes domain heterogeneity filter, w/o KD removes keyword diversity filter, w/o CKU removes complexity of knowledge units filter; (2) Environment: w/o Online uses local dense dense retrieval to synthesize training data; (3) Response Curation: w/o FR removes format regularization filter, w/o RPC removes reasoning path control filter, w/o QD removes question difficulty filter, w/o SC search count filter. As shown in Table 2, all ablated variants exhibit a decline in performance compared to our full method, underscoring the integral contribution of each component. Among them, w/o QD leads to the most significant performance drop, suggesting that question difficulty plays a crucial role in training. More challenging questions are more likely to stimulate the model’s autonomous retrieval capabilities during reasoning.
@@ 4.2 Effect of Post-SFT RL
We apply DPO and REINFORCE++ to conduct offline and online reinforcement learning, respectively. As shown in Table 3, the model trained with DPO achieves further improvements over the SFT baseline, demonstrating the effectiveness of offline preference optimization. In contrast, the model trained with REINFORCE++ produces significantly shorter responses (see Appendix C) and shows notable performance degradation on both the Bamboogle and GAIA benchmarks. This suggests that online RL mainly triggers retrieval behavior, but brings little benefit to models that are already good at retrieval. We hypothesize that the success of offline DPO stems from its ability to leverage high-quality trajectories generated by a strong LLM. These trajectories provide informative preference signals and stable supervision, allowing the model to refine its reasoning and search strategies.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) systems have advanced large language models (LLMs) in complex deep search scenarios requiring multi-step reasoning and iterative information retrieval. However, existing approaches face critical limitations that lack high-quality training trajectories or suffer from the distributional mismatches in simulated environments and prohibitive computational costs for real-world deployment. This paper introduces SimpleDeepSearcher, a lightweight yet effective framework that bridges this gap through strategic data engineering rather than complex training paradigms. Our approach synthesizes high-quality training data by simulating realistic user interactions in live web search environments, coupled with a multi-criteria curation strategy that optimizes the diversity and quality of input and output side. Experiments on five benchmarks across diverse domains demonstrate that SFT on only 871 curated samples yields significant improvements over RL-based baselines. Our work establishes SFT as a viable pathway by systematically addressing the data-scarce bottleneck, offering practical insights for efficient deep search systems. Our code and data are available at https://github.com/RUCAIBox/SimpleDeepSearcher.
@@
Shuang Sun1thanks:    Equal contributions., Huatong Song [^1] footnotemark: 1, Yuhao Wang1, Ruiyang Ren1, Jinhao Jiang1, Junjie Zhang1, Fei Bai1, Jia Deng1, Wayne Xin Zhao1thanks:    Corresponding authors., Zheng Liu2, Lei Fang322footnotemark: 2, Zhongyuan Wang2, Ji-Rong Wen1 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence 3DataCanvas Alaya NeW {sunshuanguns, batmanfly}@gmail.com
@@ 1 Introduction
In recent years, retrieval-augmented generation (RAG) methods have significantly enhanced LLMs by incorporating external knowledge retrieval mechanisms (Lewis et al., 2020; Zhao et al., 2024; Gao et al., 2024). Recent advancements have extended these capabilities to complex deep search scenarios that demand multi-step reasoning with iterative information retrieval and synthesis (Alzubi et al., 2025). Traditional RAG systems typically treat retrieval as an external auxiliary module, following a fixed pipeline of “retrieval–re-ranking–reading” (Qi et al., 2020). In contrast, deep search scenarios require the model to internalize the abilities of “when to retrieve, how to retrieve, and how to reason based on retrieved content,” in order to address more flexible and complex tasks.
To address the complex reasoning demands in deep search scenarios, early research explored prompt-based strategies that guide models to decompose questions, generate queries, and retrieve information iteratively (Jiang et al., 2024; Teng et al., 2025; Li et al., 2025a). Other studies have attempted to improve model performance through supervised fine-tuning (SFT) (Wang et al., 2025); however, there is currently a lack of high-quality trajectory data of reasoning and search interactions for training (Jin et al., 2025). To further enhance the model’s autonomous search capabilities, Reinforcement Learning (RL) (Sutton et al., 1999) is considered as a promising solution to train models through real-time interaction with the environment (Nakano et al., 2021; Song et al., 2025; Jin et al., 2025; Zheng et al., 2025). However, most RL-based approaches operate within artificial environments using static document corpora, creating a distributional mismatch with real-world web dynamics. Moreover, the inherent computational intensity of RL training escalates exponentially when interfacing with live search APIs (Sun et al., 2025).
Given the overhead and complexity of RL-based training, we hypothesize that SFT remains a viable pathway for building efficient deep search systems. While SFT offers a streamlined training process, it faces the critical challenge of lacking high-quality training data in deep search scenarios. On the one hand, existing QA datasets often lack the diversity and complexity of questions and search-oriented purposes on the Web, which are essential for deep search training. On the other hand, traditional answer annotations omit the critical reasoning traces (search operations, evidence synthesis, and efficient decision paths) required for teaching search-integrated reasoning strategies.
@@ 2 Method
Figure 1: Overall framework of our proposed SimpleDeepSearcher approach. $r$ denotes the reasoning content, $q$ represents the search query, and $d$ refers to the retrieved document after summarization. $t_{s}$ and $t_{e}$ are special tokens indicating the beginning and end of the search query, and $a$ denotes the final answer.
@@ 2.1 Overview
First, we replace static document retrieval (Karpukhin et al., 2020) with real-time network interactions, simulating human search behavior through an iterative cycle of "reasoning-searching-summarizing-generating." By directly processing raw HTML content via commercial search APIs, we capture diverse web information features—ranging from structured data snippets to unstructured narrative discourse. Based on this, we first filter input queries using domain heterogeneity, keyword diversity, and knowledge unit complexity to construct a maximally informative training foundation while ensuring selected queries align with real-world web search scenarios. Additionally, we apply a filtering mechanism to LLM-synthesized responses, implementing a four-dimensional quality filter that simultaneously optimizes format standardization, reasoning path control, question difficulty, and search effectiveness to guarantee response quality.
The framework’s modular design offers three distinctive advantages: First, it exposes the model to authentic search artifacts and noise patterns through real web interactions. Second, our multidimensional filtering strategy enables state-of-the-art performance with remarkably small SFT datasets, eliminating dependency on resource-heavy RL training. Third, the decoupled architecture between data synthesis and model constraints provides exceptional flexibility that our curated datasets can enhance any LLMs while maintaining compatibility with emerging reasoning architectures and alternative training paradigms including RL. Since the searched content is not generated by the LLM, we mask out these tokens during the SFT process.
@@ 2.2 Data Synthesis in Real Web Environment
Typically, traditional retrieval-augmented generation (RAG) systems rely on closed and static knowledge corpora (e.g., a Wikipedia snapshot). Such knowledge corpora exhibit two primary limitations: firstly, the content they contain often consists of refined and condensed segments (Chen et al., 2024); secondly, the information within these knowledge corpora lacks timeliness. Consequently, RAG systems are limited in their ability to simulate authentic user search behaviors, as users typically search within open, dynamic, and complex web environments where the information is not only diverse in format and varied in quality but is also frequently accompanied by redundancy and noise. In light of this, our data synthesis approach does not rely on curated document collections but is instead grounded in the real, open web environment. This authentic web environment also places greater demands on the model’s capabilities for information extraction, synthesis, and reasoning.
@@
Our data synthesis strategy is firmly rooted in real web scenarios, which substantially enriches the diversity and authenticity of training samples. Building on this, this strategy enables scaling of high-quality trajectory data of reasoning and search interactions.
@@ 2.3 Diversity-aware Query Sampling
To engineer a deep search architecture with advanced query comprehension and reasoning capabilities, we implement a strategic repurposing of open-domain question answering (QA) resources. These curated datasets offer natural language questions that inherently require multi-hop information retrieval operations, thereby exhibiting strong task alignment with the cognitive demands of deep search systems (Zheng et al., 2025). Our selection protocol combines single-hop and multi-hop QA benchmarks through principled composition, ensuring coverage of both atomic and composite reasoning paradigms.
However, empirical evidence suggests that naive dataset scaling yields diminishing returns in SFT (Zhou et al., 2023). The efficacy of such approaches fundamentally depends on the intrinsic diversity and informational entropy of training instances. While existing open-domain QA corpora provide substantial volume, systematic analysis reveals three critical limitations: (1) domain-specific overrepresentation creating skewed knowledge distributions, (2) repetitive syntactic patterns reducing linguistic variability (Parmar et al., 2022), and (3) semantic simplicity thresholds below real-world query complexity. These factors collectively induce model brittleness and constrain cross-domain generalization potential (see Appendix B for details). To address these critical limitations, we introduce a diversity-aware query sampling strategy to implement systematic data filtering through tripartite orthogonal criteria:
Keyword Diversity focuses on the distributional diversity of core semantic constituents (definition provided in Appendix B). we ensure non-redundant exposure to low-frequency conceptual entities, multi-order relational dependencies, and contextually ambiguous referential expressions. Such systematic variation compels the model to transcend superficial lexical pattern matching, instead developing reasoning architectures essential for interpreting complex entity interactions (Linzen, 2020).
@@
1:Annotated dataset $D$ with domains $d_{1},d_{2},\dots,d_{m}$, target number of queries NN
[…]
@@
9:   $K\leftarrow\emptyset$ $\triangleright$ Initialize the keyword set
@@ 2.4 Multi-Dimention Response Curation
The presence of low-quality reasoning outputs in language models not only compromises performance and transparency but also introduces noise into training signals, leading to inefficient computational resource utilization. To address these challenges, we developed a systematic filtering protocol that selects optimal solutions through rigorous evaluation of multiple responses per query.
@@
Model Xbench-DeepSearch BrowseComp-ZH BrowseComp-EN
Webthink-RL 24.0 7.3 2.8
WebDancer-32B 38.7 14.1 2.5
SimpleDeepSearcher 30.0 14.5 4.3
Table 2: Results on the more challenging Xbench-DeepSearch, BrowserComp-ZH, and BrowseComp-EN benchmarks. The results are evaluated with LLM-as-Judge. Results marked with * are cited from other papers or reports. The best results are in bold.
@@ 3.1 Experimental Setup
We sample training data from single-hop and multi-hop knowledge-intensive QA datasets to cover a wide range of domains and question difficulty. For single-hop questions, we use Natural Questions (Kwiatkowski et al., 2019) and SimpleQA (Wei et al., 2024). For multi-hop questions, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), MuSiQue (Tang and Yang, 2024), and MultiHopRAG (Tang and Yang, 2024). To test the model’s performance on out-of-domain data, we select Bamboogle (Press et al., 2022), FRAMES (Krishna et al., 2024), and GAIA (Mialon et al., 2023) benchmarks. In addition, we further conduct evaluations on the more challenging benchmarks including xbench-DeepSearch, BrowseComp-ZH (Zhou et al., 2025), and BrowseComp-EN (Wei et al., 2025). These datasets are not used during training and help evaluate how well the model works on new domains. We evaluate on 500 randomly sampled instances from the validation sets of HotpotQA, 2WikiMultiHopQA, and MuSiQue. For GAIA, we use 103 examples from the text-only validation subset (Li et al., 2025b), while for BrowseComp-EN we randomly sample 300 instances. For the remaining benchmarks, we use their full test sets.
@@
We consider following type of baselines: Naive Generation: Direct generation of answers without retrieval. Standard RAG(Zhao et al., 2024): Directly retrieve relevant documents by querying the original question. Search-o1 (Li et al., 2025a): Encourages the model to perform self-initiated retrieval using prompts. RAG-RL: R1-Searcher (Song et al., 2025), DeepResearcher (Zheng et al., 2025), WebThinker-RL (Li et al., 2025b), and WebDancer (Wu et al., 2025), the open-source models trained with reinforcement learning to enable self-initiated retrieval. We conduct experiments using the following model backbones with an online search engine, including Qwen-2.5-7B-Instruct, Qwen-2.5-32B-Instruct, Deepseek-Distilled-Qwen-2.5-32B, and QwQ-32B.
@@
Our experimental setup consists of three main components: SFT, generation, and query sampling. In the SFT phase, we use a total batch size of 64 and train for 6 epochs with a learning rate of 1e‑5, warmup ratio of 0.03, and a sequence length of 30,000 tokens. During fine-tuning, external retrieval documents are masked to avoid learning from noisy or spurious information. For generation, all models are configured with a maximum sequence length of 20,480 tokens, temperature of 0.6, top‑p of 0.95, and top‑k of 40. During query sampling, we used QwQ-32B to annotate each query with its corresponding domain and keywords. For data synthesis, we employed QwQ-32B as the reasoning model and Google Search API as the search engine, with a maximum of 10 search calls and 15 reasoning turns per query. For each query, we sampled 10 candidate responses. All prompts used in the experiments are provided in Appendix F.
@@ 3.2 Main Results
In addition, table 2 presents the experimental results on more complex QA datasets. These datasets are specifically designed for AI agents, requiring models to possess end-to-end planning, search, reasoning, and summarization capabilities. Notably, our model still demonstrates strong performance compared to models trained with reinforcement learning. This result further underscores the robust generalization ability of our model.
@@ 4.1 Ablation Study
To validate the effectiveness of the proposed SimpleDeepSearcher, we conduct a comprehensive ablation analysis using Qwen2.5-7B-Instruct on the Bamboogle and GAIA datasets. We conduct detailed ablation studies on three main aspects: (1) Query Sampling: w/o DD removes domain diversity filter, w/o KD removes keyword diversity filter, w/o CIW removes coverage of interrogative words filter; (2) Environment: w/o Online uses local dense dense retrieval to synthesize training data; (3) Response Curation: w/o FR removes format regularization filter, w/o RPC removes reasoning path control filter, w/o QD removes question difficulty filter, w/o SC search count filter. As observed, all ablated variants exhibit a decline in performance compared to our full method, underscoring the integral contribution of each component. Among them, w/o QD leads to the most significant performance drop, suggesting that question difficulty plays a crucial role in training. More challenging questions are more likely to stimulate the model’s autonomous retrieval capabilities during reasoning.
@@ 4.2 Effect of Post-SFT RL
We apply DPO and REINFORCE++ to conduct offline and online reinforcement learning, respectively. As shown in Table 4, the model trained with DPO achieves further improvements over the SFT baseline, demonstrating the effectiveness of offline preference optimization (see Appendix C for details). In contrast, the model trained with REINFORCE++ produces significantly shorter responses (see Appendix D for details) and shows notable performance degradation on both the Bamboogle and GAIA benchmarks. This suggests that online RL mainly triggers retrieval behavior, but brings little benefit to models that are already good at retrieval. We hypothesize that the success of offline DPO stems from its ability to leverage high-quality trajectories generated by a strong LLM. These trajectories provide informative preference signals and stable supervision, allowing the model to refine its reasoning and search strategies.